OpenStudy (anonymous):

determine whether the series converges or diverges: (-1)^n n^2/(5^n)

OpenStudy (anonymous):

\[(-1)^n \frac{ n^2 }{ 5^n }\]

OpenStudy (anonymous):

i believe it's divergent

OpenStudy (zehanz):

I looks convergent. To prove this, you have to prove that the absolute value of the number doesn't get larger and larger. I would do it this way:\[|(-1)^n \frac{n^2}{5^n}|=\frac{n^2}{5^n}< \frac{4^n}{5^n}=\left( \frac{4}{5} \right)^n \rightarrow 0 \] So the limit is 0. The series converges. Some would say this could even be done much simpler: it is well-known that exponential functions increase faster than power functions, so the denominator "wins" and the limit is 0. In fact, the argument in my calculations is the same: \(n^2 < 4^n\). The exponential sequence increases faster than the power sequence. I don't know what you are supposed to do, but these are two ways to approach it.

OpenStudy (experimentx):

try using ratio test

OpenStudy (experimentx):

there are couple of other test ... the above mentioned test is somewhat close to Leibniz test.

OpenStudy (anonymous):

it works!

OpenStudy (anonymous):

thanks to you both! how can i award both metals?

OpenStudy (zehanz):

YW! You can't :(